OpenStudy (anonymous):

i need to find the equation of the (2,10) (1,5)

OpenStudy (anonymous):

hi, well first off you need the formula for doing this problem. the formula is \[y2-y1 \over x2-x1\]

OpenStudy (anonymous):

After that you need to insert the coordinates. this would be the equation \[10-5 \over 2-1 \] which equals \[5 \over 1 \] and hat equals 5. Good luck!

OpenStudy (anonymous):

That's only the slope, Battlehero. I think he means that he needs to find the equation of the line that passes through those points. I'm not sure what form, so I'll do all three here: Since Battlehero found the slope and you have a point, we can use point slope form: \[y-y _{1} = m(x - x_{1})\] where y1 and x1 are the y and x coordinates of one of the points, and m is the slope of the line. You can use either of the points, but I'll go with (1, 5). \[y-5 = 5(x - 1)\] So that's point-slope form. If it's asking for slope intercept form, we can do some algebra to make it that. So here's slope-intercept form: \[y = mx + b\] Let's convert it. First we have to distribute the 5. \[y - 5 = 5x - 5(1)\] \[y - 5 = 5x - 5\] Add 5 to both sides. \[y - 5 + 5 = 5x - 5 + 5\] \[y = 5x\] That's slope-intercept form. Notice that the b is missing. This is because it, which is the y-intercept, is zero. For the last one, we'll do standard form. This one's the easiest. \[Ax + By = C\] So we have to get x and y on the same side. \[y - 5x = 5x-5x\] \[y - 5x = 0\] \[-5x + y = 0\] But we're not done yet. One of the rules of standard form is that A, the coefficient of x or the number x is being multiplied by, must be a positive integer. This means it must not be negative, a fraction, or a decimal. To cancel out the negative, we can divide doth sides by -1. \[\frac{-5x + y} {-1} = \frac{0}{-1}\] If you divide an expression by -1, just reverse the signs. Zero divided by -1 is zero, so here's standard form: \[5x - y = 0\]